# Quantum-enabled optical large-baseline interferometry: applications, protocols and feasibility

### v1

@@ Abstract
Optical Very Long Baseline Interferometry offers the potential for unprecedented angular resolution in both astronomical imaging and precision measurements. Classical approaches, however, face significant limitations due to photon loss, background noise, and the requirements for dynamical delay lines over large distances. This document surveys recent developments in quantum-enabled VLBI, which aim to address these challenges using entanglement-assisted protocols, quantum memory storage, and nonlocal measurement techniques. While its application to astronomy is well known, we also examine how these techniques may be extended to geodesy – specifically, the monitoring of Earth’s rotation. Particular attention is given to quantum-enhanced telescope architectures, including repeater-based long-baseline interferometry and quantum error-corrected encoding schemes, which offer a pathway toward high-fidelity optical VLBI. To aid the discussion, we also compare specifications for key enabling technologies to current state-of-the-art experimental components, including switching rates, gate times, entanglement distribution rates, and memory lifetimes. By integrating quantum technologies, future interferometric networks may achieve diffraction-limited imaging at optical and near-infrared wavelengths, surpassing the constraints of classical techniques and enabling new precision tests of astrophysical and fundamental physics phenomena.
@@ I Introduction
Several challenges hinder the progress in building large-baseline optical interferometers, including the weight of the instrument itself, photon loss, background noise, and the requirements for long dynamical delay lines. These factors ultimately limit the achievable distance between telescope sites. Quantum technologies can help bypass transmission losses: using quantum memories and entanglement, we can replace the direct optical links, allowing for much larger distances.
Among the key scientific applications of optical Very Large Baseline Interferometry (VLBI), two stand out. First, VLBI offers a path toward high angular resolution astronomical imaging, particularly in regimes that are inaccessible to conventional telescopes. Second, and less commonly considered, optical VLBI can achieve unprecedented spatial and temporal resolution in geodetic metrology, such as monitoring Earth’s rotation, tectonic motion, and lunar libration.
Today’s most advanced interferometric imaging systems operate at radio and microwave wavelengths. Instruments such as the Event Horizon Telescope (EHT) achieve resolutions of tens of microarcseconds by combining a maximum baseline of $\sim$12,000 km (roughly the Earth’s diameter) with a typical wavelength of 3 mm. Shifting to optical wavelengths can, in principle, increase resolution by several orders of magnitude. For instance, an optical system operating at $\sim$600 nm would outperform the EHT in resolution with a baseline as short as 2.4 km. At 1550 nm (a standard C-band telecom wavelength), the equivalent threshold is 6.2 km.
@@ II Astronomical imaging
The underlying physics of the black hole imaging setup is purely classical: the array operates at radio frequencies, where both the amplitude and phase of the received electromagnetic field can be directly measured. The data is then post-processed to reconstruct an image. Moving to optical frequencies, however, introduces significant complications: even the fastest electronics cannot directly measure the oscillations of the electric field at optical frequencies. Nevertheless, if these complications can be overcome, the potential gains in resolution are substantial. Optical VLBI offers a 3-5 orders of magnitude improvement in resolution over radio-based systems, depending on the wavelength and baseline. This opens the door to imaging astrophysical structures that are beyond the reach of our current imaging systems.
Optical VLBI refers to interferometric regimes where real-time calibration of optical path lengths is not feasible, necessitating the use of closure phases and post-processed visibilities [7]. Instruments like CHARA [8] and VLTI [9] operate in this regime, and future concepts such as the Big Fringe Telescope [10] aim to extend it. Our work does not propose optical VLBI as a new idea, but rather explores how quantum-enhanced protocols might overcome key technical limitations, particularly in photon loss and scalability.
@@
Current optical interferometers face severe constraints on scalability and baseline length, due to the need for physical delay lines and lossy multi-way beam combination. Quantum architectures offer a route to nonlocal beam combination by connecting telescopes without direct optical interference. If deployed on the Moon, we can exploit the natural vacuum as a beam guide for long-baseline operation without atmospheric distortion.
In the remainder of this section discuss some science cases for using large-baseline optical interferometry, including imaging of exoplanets, tracking the stars around black holes, and measuring separations of double-star systems.
@@ II.0.1 Imaging exoplanets
Exoplanets present a formidable observational challenge: they are both extremely small in angular size and orders of magnitude fainter than their host stars. Direct imaging of these objects requires two key capabilities:
  • High angular resolution, required to spatially separate the planet from the glare of the star.
  • High visibility (contrast), required to detect the planet’s faint signal against the much brighter stellar companion.
Optical VLBI offers a potential solution to both challenges. At a wavelength of 600 nm and a baseline of 100 km, the achievable angular resolution is approximately 1.24 microarcseconds ($\upmu$as)—well beyond the reach of existing instruments, including the Event Horizon Telescope. At this resolution, it becomes conceivable to directly image spatially resolved features in the atmospheres of nearby exoplanets, opening new possibilities for studying their composition, structure, and potentially even weather systems. We show the separation of the star from the exoplanet in Table 1.
@@ II.0.2 Tracking stars orbiting black holes
where $\chi$ is the dimensionless spin parameter of the black hole (ranging from 0 to 1), and $R_{S}$ is the Schwarzschild radius. For the star S2, which orbits the supermassive black hole at the centre of our galaxy, the Lense–Thirring precession is of the order of 0.05 arcminutes per orbit — about 240 times smaller than the Schwarzschild precession [13]. Resolving such small effects requires extremely high angular resolution. The GRAVITY collaboration (see, e.g., Refs. [14, 15, 16]) currently achieves near-infrared resolution of 3 milli-arcseconds and astrometric precision of 10–100 $\upmu$as, and a baseline of 130 m. To detect the Lense–Thirring precession at comparable wavelengths (near infrared), a baseline of at least $130\times 240=32$ km would be required.
Figure 1: The Lense–Thirring effect: a star’s orbit precesses due to the frame-dragging caused by a rotating massive object, such as a spinning black hole.
@@ III Geodesy
Figure 2: Geometry for using VLBI for geodesy. The position of the reference star is assumed fixed, and the goal is to measure the baseline vector $\bm{B}$.
VLBI is a geometric technique; it measures the time difference between the arrival at two Earth-based antennas emitted by a distant quasar. The basic geometric delay $\tau$ measured by two radio telescopes with Earth-based positions $r_{1}$ and $r_{2}$ is given by
@@
The core idea is the precise timing of light signals:
  1. 1.
    When a quasar emits EM waves, they arrive at different telescopes at slightly different times due to Earth’s shape and rotation.
  2. 2.
    The exact time delay between telescopes depends on the Earth’s rotation state at that moment.
  3. 3.
    By comparing the signals and modelling their delay based on Earth’s motion, scientists can infer tiny variations in rotation speed and axis position.
Geodetic VLBI plays a crucial role in geodesy, enabling precise measurements of the Earth’s radio telescope positions and long-term motion of the tectonic plates. Currently, the tectonic motion is measured with accuracy up to 0.1 mm/year, though the seasonal variations of the individual radio telescopes’ positions are also monitored [17]. Some of the radio telescopes were affected by strong earthquakes (Japan, 2010; Chile, 2011) [18]. Hence, the co-seismic and post-seismic displacement can be monitored. A global catastrophe, like the Sumatra Boxing Day earthquake on 26-Dec-2004, may cause a global shift in positions of all geodetic sites around the Earth with an amplitude of $\sim$1 mm. In addition, geodetic VLBI studies the post-glacial uplift in Fenno-Scandinavia, the northern part of North America, and Antarctica. The global melting of glaciers causes a dramatic change in the immediate positions of radio telescopes, and changes the shape of the planet due to the secular shift of the inertia momentum.
@@ III.1.1 Time Scale
The leap second problem arises from the need to synchronise two fundamentally different timekeeping systems: Atomic Time (TAI/UTC) and Astronomical Time (UT1). Atomic Time, based on hydrogen and caesium atomic clocks, is highly stable and defines the SI second with unparalleled precision. In contrast, Astronomical Time (UT1) is tied to Earth’s rotation, which varies unpredictably. As a result, UT1 drifts relative to atomic time, necessitating the periodic introduction of leap seconds to maintain synchronisation. In the 1990s, the leap second was introduced almost every year. Although the Earth rotation deceleration nowadays is not so high and the leap second is introduced rarely (the last time was in 2015), the situation may change again and needs to be monitored continually [19].
Leap seconds introduce significant challenges for computing, satellite navigation, and global timekeeping systems. Since leap seconds occur irregularly and unpredictably, they disrupt time-dependent operations, requiring software updates and manual interventions. To mitigate these issues, discussions are ongoing regarding the potential elimination or redefinition of leap seconds in the future [20]. Optical VLBI offers substantial improvement in the temporal resolution of UT1 measurements, enabling detection of changes in Earth’s rotation on shorter and more precise timescales.
@@ III.1.2 Polar motion
Polar motion is the motion of the instantaneous Earth rotation axis around the geographic pole. The regular motion consists of two major components, namely the annual rotation with a period of approximately 365 days and the Chandler wobble with a period of approximately 430 days. The Chandler wobble is a free rotation effect with variable period and amplitude. Currently, it has disappeared from the total Earth’s pole motion (presumably temporarily) for unknown reasons. VLBI helps in tracking polar motion, as the Earth’s rotational axis drifts slightly due to mass redistributions, such as the melting of ice sheets and shifts in the planet’s mantle [21][22]. Optical VLBI can enhance our ability to track polar motion with finer spatial precision due to the shorter wavelengths involved, potentially improving sensitivity to small-scale mass redistributions.
@@ III.2 Astrometry
The main ICRF3 reference catalogue is based on observations at a frequency of 8.4 GHz (the X band). In parallel, there are additional catalogues for frequencies of 24 GHz and 32 GHz with fewer objects. These catalogues are important for tracking deep space missions [28]. If optical VLBI is used, the enhancement is a finer determination of source positions in the celestial reference frame, with potential uncertainties reduced by 3-5 orders of magnitude compared to radio VLBI.
@@ IV The model and parameter estimation
High-resolution astronomical imaging and high-precision geodesy are fundamentally linked by their reliance on the same core technology: interferometry. In astronomical imaging, interferometers are used to resolve fine spatial structure by coherently combining light collected from a distant source. In geodesy, the principle is inverted: rather than using interferometry to study a celestial target, the celestial source serves as a fixed reference, and the interferometric measurement is used to track the motion of the Earth itself. In both contexts, the required technology is essentially the same, therefore, in this work, we adopt a common physical model for the stellar source. To facilitate analytic comparison and benchmarking, we focus on the well-understood two-mode model, for which the quantum Fisher information is readily computable and serves as a natural metric of the resolution.
@@
where we have used the quadrature ordering $(q_{A},p_{A},q_{B},p_{B})$, with subscripts referring to Alice ($A$) and Bob ($B$).
In the limit that $\epsilon\ll 1$, which is typically true for astronomy, we can describe the optical state by the density matrix,
@@
If we were to divide a bandwidth spectrum (centred at 500 THz or 600 nm) of $\Delta\lambda=10$ nm into 100 MHz narrow bands, this is approximately 8300 bands; the photon will have a coherence time of approximately 10 ns with a coherence length 3 m; if instead we divide this into 1 GHz bands (830 bands), then the coherence time will be 1 ns, and the coherence length will be 30 cm. In the literature, the bandwidth of the stellar source is often quoted as either frequency or wavelength; we include Table 4 for reference and ease of converting between them.
@@
where $N$ is the number of copies of $\hat{\rho}_{\theta}$ used and $J_{\theta}$ is the quantum Fisher information (QFI) associated with the state $\hat{\rho}_{\theta}$.
If there are multiple parameters we want to estimate, where $\bm{\mathrm{\theta}}=(\theta_{1},\theta_{2},\dots)$, we can define a QFI matrix $\bm{\mathrm{J}}$ that quantifies not only the QFI for each parameter (diagonal components) but also for correlated parameters (off-diagonal components). The matrix elements are given by
@@
where $\hat{L}_{j}$ is the symmetric logarithmic derivative with respect to $\theta_{j}$ [35].
The inverse of the QFI matrix provides a lower bound on the covariance matrix $\big{[}\operatorname{Cov}(\bm{\mathrm{\theta}})\big{]}_{jk}=\braket{\theta_{j}\theta_{k}}-\braket{\theta_{j}}\braket{\theta_{k}}$,
@@
The QFI matrix elements for the incoming stellar state are found to be
@@
$J_{\phi\gamma}$ $=0$ (15c)
If we take the equivalent trace norm of the QFI matrix,
@@
Note that the above expression can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$. It’s also perhaps unfair to lump them together because $J_{\phi}/\epsilon$ is capped by $1$, whereas $J_{\gamma}/\epsilon$ can, in principle, approach infinity.
@@
      Requirements       Direct detection       quantum-enabled
@@
$\phi$ $=\omega\tau,$
$=\frac{B\sin\theta}{\lambda}.$ (17)
From quantum metrology, we know that the QFI of $\phi$ is equal to 1; i.e. $\Delta\phi\approx 1$. Here we assume $\theta$ to be fixed, and $\lambda$ is also fixed. This means
@@
The relation in Eq. (18) reveals a key advantage of operating at optical wavelengths: since the uncertainty in the baseline measurement scales inversely with $\lambda$, moving from radio to optical can potentially yield a substantial gain in precision. For the same number of detected photons and observational geometry, shorter wavelengths directly translate into increased precision of geodetic measurements.
V Technological requirements of OVLBI

### v2

@@ Abstract
Optical Very Long Baseline Interferometry (VLBI) offers the potential for unprecedented angular resolution in both astronomical imaging and geodesy measurements. Classical approaches face limitations due to photon loss, background noise, and their need for dynamical delay lines over large distances. This document surveys recent developments in quantum-enabled optical VLBI, which aim to address these challenges using entanglement-assisted protocols, quantum memory storage, and nonlocal measurement techniques. While its application to astronomy is well known, we also examine how these techniques may be extended to geodesy – specifically, the monitoring of Earth’s rotation. Particular attention is given to quantum-enhanced telescope architectures, including repeater-based long-baseline interferometry and quantum error-corrected encoding schemes, which offer a pathway toward high-fidelity optical VLBI. To aid the discussion, we also compare specifications for key enabling technologies to current state-of-the-art experimental components. By integrating quantum technologies, future interferometric networks may achieve diffraction-limited imaging at optical and near-infrared wavelengths, surpassing the constraints of classical techniques and enabling new precision tests of astrophysical and fundamental physics phenomena.
@@ I Introduction
The underlying physics of the black hole imaging setup is purely classical: the array operates at radio frequencies, where both the amplitude and phase of the received electromagnetic field can be directly measured. The data is then post-processed to reconstruct an image. However, direct measurement is not possible in optics, because even the fastest electronics cannot directly measure the oscillations of the electric field at optical frequencies. Nevertheless, if we can find alternate solutions, we can potentially gain 3-5 orders of magnitude improvement in resolution. This unlocks possibilities for imaging astrophysical structures that are beyond the reach of our current imaging systems.
Several challenges hinder the progress in building large-baseline optical interferometers, including the weight of the instrument itself, photon loss, background noise, and the requirements for long dynamical delay lines. These factors ultimately limit the achievable distance between telescope sites. Quantum technologies can help bypass transmission losses: using quantum memories and entanglement, we can replace the direct optical links, allowing for much larger distances. If deployed on the Moon, we can exploit the natural vacuum as a beam guide for long-baseline operation without atmospheric distortion.
Among the key scientific applications of optical VLBI, two stand out. First, VLBI offers a path toward high angular resolution astronomical imaging. Second, and less commonly considered, optical VLBI can achieve unprecedented spatial and temporal resolution in geodetic metrology, such as monitoring Earth’s rotation, tectonic motion, and lunar libration.
Optical VLBI refers to interferometric regimes where real-time compensation of optical path lengths is not feasible, necessitating the use of closure phases and post-processed visibilities [7]. Instruments like CHARA [8] and VLTI [9] operate in this regime, and future concepts such as the Big Fringe Telescope [10] aim to extend it. Our work does not propose optical VLBI as a new idea, but rather explores how quantum-enabled protocols might overcome key technical limitations, particularly in photon loss and scalability.
Today’s most advanced interferometric imaging systems operate at radio and microwave wavelengths. Instruments such as the Event Horizon Telescope (EHT) achieve resolutions of tens of microarcseconds by combining a maximum baseline of $\sim$12,000 km (roughly the Earth’s diameter) with a typical wavelength of 3 mm. Shifting to optical wavelengths can potentially increase resolution by several orders of magnitude. An optical system operating at 600 nm could match the EHT with only a 2.4 km baseline; at 1550 nm, 6.2 km would suffice.
@@ II Astronomical imaging
This section discusses some science cases for using large-baseline optical interferometry, including imaging of exoplanets, tracking the stars around black holes, and measuring separations of double-star systems.
@@ II.0.1 Imaging exoplanets
Imaging exoplanets is a challenging task: they are both extremely small in angular size and orders of magnitude fainter than their host stars. Direct imaging of these objects requires:
  • high angular resolution to spatially separate the planet from the glare of the star.
  • high visibility (contrast) to detect the planet’s faint signal against the much brighter parent star.
Optical VLBI offers a potential solution to both challenges. At a wavelength of 600 nm and a baseline of 100 km, the achievable angular resolution is approximately 1.24 microarcseconds ($\upmu$as)—well beyond the reach of existing instruments, including the Event Horizon Telescope. At this resolution, it becomes conceivable to directly image spatially resolved features in the atmospheres of nearby exoplanets, opening new possibilities for studying their composition, structure, and potentially even weather systems. We show the separation of the star from the exoplanet in Table 1 for reference.
@@ II.0.2 Tracking stars orbiting black holes
where $\chi$ is the dimensionless spin parameter of the black hole (ranging from 0 to 1), and $R_{S}$ is the Schwarzschild radius. For the star S2, which orbits the supermassive black hole at the centre of our galaxy, the Lense–Thirring precession is of the order of 0.05 arcminutes per orbit — about 240 times smaller than the Schwarzschild precession [13]. Resolving such small effects requires extremely high angular resolution. The GRAVITY collaboration (see, e.g., Refs. [14, 15, 16]) currently achieves a near-infrared resolution of 3 milli-arcseconds and astrometric precision of 10–100 $\upmu$as, and a baseline of 130 m. To detect the Lense–Thirring precession at comparable wavelengths (near infrared), a baseline of at least $130\times 240=32$ km would be required.
Figure 1: The Thirring effect: a star’s orbit precesses due to the frame-dragging caused by a rotating massive object, such as a spinning black hole.
@@ III Geodesy
Figure 2: Geometry for using VLBI for geodesy. The position of the reference star is assumed to be fixed, and the goal is to measure the baseline vector $\bm{B}$.
Geodesy uses precise measurements of Earth’s shape, orientation, and gravity field—often through VLBI and satellite techniques—to monitor tectonic motion, the Earth’s rotation, and maintain global reference frames. VLBI is a geometric technique; it measures the time difference between the arrival at two Earth-based antennas emitted by a distant quasar. The basic geometric delay $\tau$ measured by two radio telescopes with Earth-based positions $r_{1}$ and $r_{2}$ is given by
@@
When a quasar emits EM waves, they arrive at different telescopes at slightly different times due to Earth’s shape and rotation. The exact time delay between telescopes depends on the Earth’s rotation state at that moment. By comparing the signals and modelling their delay based on Earth’s motion, we can infer small variations in rotation speed and axis position.
Geodetic VLBI plays a crucial role in geodesy. Currently, the tectonic motion is measured with an accuracy up to 0.1 mm/year, though the seasonal variations of the individual radio telescopes’ positions are also monitored [17]. Some of the radio telescopes were affected by strong earthquakes (Japan, 2010; Chile, 2011) [18]. Hence, the co-seismic and post-seismic displacement can be monitored. A global catastrophe, like the Sumatra Boxing Day earthquake on 26-Dec-2004, may cause a global shift in positions of all geodetic sites around the Earth with an amplitude of $\sim$1 mm. In addition, geodetic VLBI studies the post-glacial uplift in Fenno-Scandinavia, the northern part of North America, and Antarctica. The global melting of glaciers causes a dramatic change in the immediate positions of radio telescopes, and changes the shape of the planet due to the secular shift of the inertia momentum.
@@ III.1.1 Time Scale
The leap second problem arises from the need to synchronise two fundamentally different timekeeping systems: Atomic Time (TAI/UTC) and Astronomical Time (UT1). Atomic Time, based on hydrogen and caesium atomic clocks, is highly stable and defines the SI second with unparalleled precision. In contrast, Astronomical Time (UT1) is tied to Earth’s rotation, which varies unpredictably. As a result, UT1 drifts relative to atomic time, necessitating the periodic introduction of leap seconds to maintain synchronisation. In the 1990s, the leap second was introduced almost every year. Although the Earth’s rotation deceleration nowadays is not so high and the leap second is introduced rarely (the last time was in 2015), the situation may change again and needs to be monitored continually [19].
Leap seconds introduce significant challenges for computing, satellite navigation, and global timekeeping systems. Since leap seconds occur irregularly and unpredictably, they disrupt time-dependent operations, requiring software updates and manual interventions. To mitigate these issues, discussions are ongoing regarding the potential elimination or redefinition of leap seconds in the future [20]. Optical VLBI offers a substantial improvement in the temporal resolution of UT1 measurements, enabling the detection of changes in Earth’s rotation on shorter and more precise timescales.
@@ III.1.2 Polar motion
Polar motion is the motion of the instantaneous Earth rotation axis around the geographic pole. The regular motion consists of two major components, namely the annual rotation with a period of approximately 365 days and the Chandler wobble with a period of approximately 430 days. The Chandler wobble is a free rotation effect with variable period and amplitude. Currently, it has disappeared from the total Earth’s pole motion (presumably temporarily) for unknown reasons. VLBI helps in tracking polar motion, as the Earth’s rotational axis drifts slightly due to mass redistributions, such as the melting of ice sheets and shifts in the planet’s mantle [21][22]. Optical VLBI can enhance our ability to track polar motion with finer spatial resolution, potentially improving sensitivity to small-scale mass redistributions.
@@ III.2 Astrometry
The main ICRF3 reference catalogue is based on observations at a frequency of 8.4 GHz (the X band). In parallel, there are additional catalogues for frequencies of 24 GHz and 32 GHz with fewer objects. These catalogues are important for tracking deep space missions [28].
@@ IV The model and parameter estimation
High-resolution astronomical imaging and high-precision geodesy are fundamentally linked by their reliance on the same core technology: interferometry. In astronomical imaging, interferometers are used to resolve fine spatial structures by coherently combining light collected from a distant source. In geodesy, the principle is inverted: the celestial source serves as a fixed reference, and the interferometric measurement is used to track the motion of the Earth itself. Given the shared technology, we adopt a common physical model for the stellar source and focus on the well-understood two-mode case, where quantum Fisher information serves as a natural metric for comparison.
@@
where we have used the quadrature ordering $(q_{A},p_{A},q_{B},p_{B})$, with subscripts referring to Alice ($A$) and Bob ($B$). In the limit that $\epsilon\ll 1$, which is typically true for astronomy, we can describe the state by the density matrix,
@@
$N$ is the number of copies of $\hat{\rho}_{\theta}$ used and $J_{\theta}$ is the quantum Fisher information (QFI) associated with the state $\hat{\rho}_{\theta}$.
If there are multiple parameters we want to estimate, where $\bm{\mathrm{\theta}}=(\theta_{1},\theta_{2},\dots)$, we can define a QFI matrix $\bm{\mathrm{J}}$. The matrix elements are given by
@@
where $\hat{L}_{j}$ is the symmetric logarithmic derivative with respect to $\theta_{j}$ [35]. The inverse of the QFI matrix provides a lower bound on the covariance matrix $\big{[}\operatorname{Cov}(\bm{\mathrm{\theta}})\big{]}_{jk}=\braket{\theta_{j}\theta_{k}}-\braket{\theta_{j}}\braket{\theta_{k}}$,
@@
The QFI matrix elements for the incoming stellar state are [36]
@@
$J_{\phi\gamma}$ $=0.$ (15c)
If we take the trace norm of the QFI matrix,
@@
Note that Eq. (IV) can be significantly larger than $\epsilon$, and that care should be taken to discuss the sum $J_{\phi}$ of $J_{\gamma}$, because because $J_{\phi}/\epsilon\leq 1$, whereas $J_{\gamma}/\epsilon$ can, in principle, approach infinity.
@@
      Requirements       Direct interferometry       quantum-enabled
@@
$\phi$ $=\omega\tau=\frac{B\sin\theta}{\lambda}.$ (17)
From quantum metrology theory, we know that the QFI of $\phi$ is equal to 1; i.e. $\Delta\phi\approx 1$. Assume $\theta$ and $\lambda$ to be fixed, this means
@@
Eq. (18) highlights a key advantage of optical wavelengths: because baseline uncertainty scales inversely with $\lambda$, shorter wavelengths significantly improve geodetic precision for the same photon count and geometry.
V Technological requirements of Optical VLBI